OpenStudy (saad560):

1.4x−0.4x+1.8=2.6.What is the value of x? -.2 .8 4.4 5.4

OpenStudy (saad560):

@triciaal

OpenStudy (saad560):

@GIitter

563blackghost (563blackghost):

Combine like terms.... \(\huge{1.4x - 0.4x=?}\)

OpenStudy (saad560):

1

OpenStudy (triciaal):

follow the steps all the terms with the unknown to one side and all the constants to the other. keep the balance whatever you do to one sideof the equation also do to the other simplify divide each side by the coefficient of the unknown variable

563blackghost (563blackghost):

Correct so that would simplify to `x`.... \(\huge{x+1.8=2.6}\) Next we need to isolate x so we would subtract 1.8 from both sides... \(\huge{x+(1.8-1.8)=2.6-1.8}\)

OpenStudy (saad560):

.8

563blackghost (563blackghost):

\(\Huge{\checkmark}\)